Circular No. 6481

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 960720 T. Murakami, R. Shibata, H. Inoue, and F. Nagase, Institute of Space

and Astronautical Science; A. Yoshida, N. Kawai, C. Otani, and F. Tokanai, Institute of Physical and Chemical Research; and J. Heise, Space Research Organization of the Netherlands, Utrecht, report on behalf of the ASCA team: “We observed the error region of the SAX-BATSE γ-ray burst (IAUC 6467) with ASCA on Sept. 19 with a net exposure of 26 000 s. Only one weak x-ray source was detected within the SAX-WFC error region, whose location is consistent with the reported x-ray source of the ROSAT all-sky survey (IAUC 6469), to an accuracy of within 1′ radius. The observed GIS countrate (0.7–10 keV) was very weak, (2.8 ± 0.6) × 10^{−3} count/s. Assuming a power-law model with index −2, and a galactic absorption toward this direction of 2.5 × 10^{20} cm^{−2}, the observed flux was 1 × 10^{−13} erg cm^{−2} s^{−1} in the band 0.5–2 keV and 2 × 10^{−13} erg cm^{−2} s^{−1} in the band 0.5–10 keV. The flux at 0.5–2 keV is roughly consistent with the flux observed with ROSAT before the burst that occurred during 1990 Aug. 22-28 (IAUC 6469).”

Corrigenda to IAUC 6480. Line 22, for (2.5 ± 0.5) read (0.9 ± 0.2). Line 27, for ∼ 1 × 10^{−13} read ∼ 4 × 10^{−14}. Line 28, for 1 and 3 × 10^{−13} read 5 and 10 × 10^{−14}.

COMET C/1996 E1 (NEAT) Continuation to ephemeris on IAUC 6440, from orbital elements by

B. G. Marsden on MPC 27882:

1996TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Oct. 4 19^{h}19.^{m}25 +48^{o}19.^{′}5 1.196 1.683 99.^{o}7 35.^{o}9 11.2
9 19 30.85 +40 59.8 1.242 1.725 100.1 34.7 11.3
14 19 40.63 +34 14.0 1.307 1.768  99.4 33.8 11.6
19 19 49.25 +28 10.1 1.390 1.813  97.6 33.0 11.8
24 19 57.07 +22 50.3 1.486 1.859  95.0 32.2 12.1
29 20 04.33 +18 13.1 1.594 1.906  91.9 31.4 12.3
Nov. 3 20 11.19 +14 14.4 1.711 1.953  88.4 30.5 12.6
8 20 17.74 +10 49.7 1.835 2.002  84.6 29.5 12.8
13 20 24.08 + 7 54.4 1.963 2.051  80.7 28.4 13.1
18 20 30.25 + 5 24.2 2.094 2.100  76.7 27.3 13.3
23 20 36.27 + 3 15.3 2.227 2.150  72.7 26.0 13.6
1996 October 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green